OpenStudy (anonymous):

What is the rate of change for f(x) = 4 sin x − 2 on the interval from x = 0 to x = pi over 2?

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

rate of change = change_in_Y/change_in_X

OpenStudy (anonymous):

So how would I do it using that problem?

OpenStudy (dan815):

okay whats the height at x=0?

OpenStudy (dan815):

you are given this equation f(x) = 4 sin x − 2

OpenStudy (dan815):

your change in height is change in f(x)

OpenStudy (anonymous):

Would it be -4?

OpenStudy (dan815):

no

OpenStudy (dan815):

sub in x=0 into that equation

OpenStudy (dan815):

f(x) = 4 sin x − 2 f(0)=4*sin(0)-2 =?

OpenStudy (dan815):

this is what you are basically doing on a graph,